### v1

# Schemex: Interactive Structural Abstraction from Examples with Contrastive Refinement

Abstract.
Each type of creative or communicative work is underpinned by an implicit structure. People learn these structures from examples—a process known in cognitive science as schema induction. However, inducing schemas is challenging, as structural patterns are often obscured by surface-level variation. We present Schemex, an interactive visual workflow that scaffolds schema induction through clustering, abstraction, and contrastive refinement. Schemex supports users through visual representations and interactive exploration that connect abstract structures to concrete examples, promoting transparency, adaptability, and effective human-AI collaboration. In our user study, participants reported significantly greater insight and confidence in the schemas developed with Schemex compared to those created using a baseline of an AI reasoning model. We conclude by discussing the broader implications of structural abstraction and contrastive refinement across domains.
generative AI, structural abstraction, contrastive refinement, schema induction, structure discovery
$c$s: Human-centered computing Interactive systems and tools
Figure 1. Schemex is an interactive, AI-powered visual workflow that helps users induce schemas from examples. It comprises three main stages: clustering, abstraction, and contrastive refinement.
1. Introduction
All creative or communicative works are underpinned by structural frameworks. When people undertake tasks such as writing a research abstract, designing a slide deck, or making a product demo, they are rarely starting from a blank slate. Instead, they draw on an internalized understanding of how such artifacts are typically organized—what components are expected, how they are sequenced, and what makes the result coherent and effective. This structural knowledge, while often tacit rather than explicit, plays a critical role in both interpreting existing examples and producing new ones. By examining multiple instances of a genre or task, people begin to identify recurring patterns—mental templates that organize content and guide its construction. The clearer and more explicit this structure becomes, the better it supports purposeful, flexible creation rather than mere surface-level imitation. In cognitive science, such internalized structural representations are known as schemas: abstract models that capture the essential components of a concept and the relationships among them.
Schema induction is the process of discovering the abstract structure from unstructured examples. Experts often create schemas implicitly from years of analyzing and creating examples. However, schemas can also be made explicit—these explicit schemas (also known as design patterns) are extremely helpful to novices in structuring and guiding their creative process. The cognitive process of schema induction broadly resembles the iterative process of sensemaking, incorporating both inductive and deductive approaches. One must iteratively verify that the abstractions and structures induced from the examples actually fit those examples, and when discrepancies arise, deduce how to reconcile them. For example, from looking at examples of HCI papers, one might learn that the first sentence is typically broad motivation, meant to show why the problem or areas the paper addresses are important. By looking at multiple examples, the recurring patterns—and variations within them, can be deduced.
However, inducing schemas from examples is challenging because of the vast amount of ambiguity in identifying hidden structures. First, structural patterns are often obscured by surface variation, as examples use diverse language and achieve similar goals by different means. For example, motivation sentences in abstracts all motivate different topics and in different ways using different words, but they all achieve the same goal. Second, examples within a set may adhere to different underlying schemas. For example, a system paper and a study paper contain different elements—while their schemas are related, they are distinct. Trying to generate a schema that explains both will likely be overly general. Third, schemas are difficult to evaluate—often the abstractions and structures in a schema can only truly be evaluated when they are applied to multiple examples to test the “fit” and generalizability. Lastly, schemas must balance between specificity and generalizability: overly abstract schemas offer insufficient guidance, while excessively specific ones transfer poorly across contexts.
To address these challenges, we frame schema induction as an interactive process grounded in real-world examples and guided by AI-assisted pattern discovery. Drawing on recent advances in AI reasoning—such as DeepSeek R1’s “slow thinking” capabilities (Guo et al., 2025)—we explore how AI can scaffold the core cognitive steps of clustering, abstraction, and contrastive refinement.
  • Clustering: Starting from a set of examples, the AI identifies latent groupings based on structural similarities and divides examples into groups.
  • Abstraction: Within each cluster, it infers underlying dimensions and derives both overall and dimension-specific attributes.
  • Contrastive Refinement: The current schema is used to generate outputs, which are then compared to real (“gold”) examples and iterate the schema. To guard against overfitting, unseen validation examples are also included.
Throughout the process, humans play a central role as reflective evaluators—they make critical judgments about coherence and utility, direct the exploration process, and retain decision agency over when to advance or refine the schema. This approach grounds schema induction in concrete examples, promoting transparency, adaptability, and effective human-AI collaboration.
To operationalize this approach, we developed Schemex, an interactive visual workflow for schema induction. Schemex scaffolds interactive structural abstraction and constrastive refinement through AI-assisted reasoning and visualizations that keep users grounded in real examples. Users initiate the process by specifying a goal (e.g., “Write HCI paper abstracts”) and providing a set of examples (e.g., CHI best paper abstracts). These examples can be multimodal; Schemex employs models such as GPT-4V and Whisper to convert images or videos into structured textual descriptions. The system clusters examples based on structural similarity and displays them as interactive visual nodes. Users can explore these clusters to inspect shared features and example-level mappings, validating the coherence of the clustering. They can then select a cluster of interest (e.g., empirical studies) and prompt the system to extract key structural dimensions (e.g., Motivation, Method, Findings). A visualized matrix allows users to examine how strongly each example reflects each dimension, supported by in-line explanations and citations. Users guide the next step by requesting inferences of dimension-level attributes (e.g., “Challenge Statement” under Motivation) and overall attributes (e.g., tone, length, or style). They evaluate the coherence of these attributes across examples and explore implementation details through contextual snippets. Users then apply the derived schema: given an input (e.g., a paper title), the system generates content guided by the schema. Schemex enables users to compare generated outputs and original examples via a color-coded dimensional analysis interface. Differences are highlighted, and the system suggests schema refinements that can be incorporated in the iteration. Validation cases not used in earlier steps are also provided, allowing further testing of schema robustness. This iterative cycle—observe, apply, compare, refine—continues until users arrive at a coherent, actionable schema tailored to their goal. In our user study, participants reported significantly greater insight and confidence in the schemas developed with Schemex compared to those created with the baseline condition, OpenAI o1-pro. Moreover, both schemas and outputs generated with Schemex were rated significantly higher in quality and depth by expert evaluators compared to those produced using the baseline condition.
The paper makes the following contributions:
  • A structural abstraction technique that leverages clustering and pattern analysis to infer dimensions and attributes from example sets, including multimodal artifacts;
  • A contrastive refinement method that supports structured generation tasks by comparing AI-generated and real examples to guide schema evolution;
  • An interactive visual workflow that links inferred schema elements to source examples, enabling interpretable and revisable schema induction;
  • An empirical evaluation showing that schemas and outputs produced with Schemex are rated significantly higher in quality and example alignment than those from a strong AI baseline (o1-pro).
2. Related Work
2.1. Schema Induction and Cognitive Foundations
Humans can infer schemas from concrete experiences. Cognitive science theories of analogy and structural alignment suggest that comparing multiple examples enables people to abstract their common relational structure (Gick and Holyoak, 1983; Gentner and Markman, 1997). In classic analogical problem-solving studies, Gick and Holyoak (Gick and Holyoak, 1983) demonstrated that examining two analogs can lead to the induction of a schema—a generalized solution framework that preserves shared relations while ignoring surface differences. Such analogical schemas support the transfer of knowledge to novel problems. These cognitive theories provide a foundation for the idea that people can induce schemas or conceptual models by observing recurring patterns across multiple instances. An effective schema abstracts away surface details to highlight underlying structure. For example, prior work on schema induction (Yu et al., 2014a) mentions that, to capture the essence of an invention, it is more useful to focus on its purpose (e.g., detaching components) and mechanism (e.g., using comb-like features), rather than vague notions (e.g., “making things easy”) or overly specific traits (e.g., color). However, striking the right balance between specificity and generalizability remains a challenge in schema induction.
Computational models in AI and cognitive science have sought to replicate this human capacity for structure learning. Kemp and Tenenbaum’s hierarchical Bayesian model identifies the most appropriate structural form (such as trees, rings, or clusters) from a predefined space to explain a dataset (Kemp and Tenenbaum, 2008). It treats structure induction as a model selection problem across graph-based representations and has successfully uncovered latent structures in domains ranging from biology to social networks. Other models emphasize the learning of relational schemas. For example, DORA, a neural-symbolic model, discovers new relational concepts through analogy and represents them as explicit predicates or schema-like structures (Doumas et al., 2008). However, such models often lack grounding in real-world artifacts, limiting their practical applicability.
Together, these cognitive and computational foundations suggest that inferring schemas through analogical generalization and structure induction is both a core human strategy and a possible computational goal. Schemex builds on this foundation by enabling users to perform schema induction interactively—leveraging human pattern recognition alongside algorithmic assistance to extract structural abstractions from example collections.
2.2. Approaches of Interactive Structural Abstraction and Contrastive Refinement
Interactive systems increasingly support users in forming abstractions from complex data by integrating direct manipulation with automated inference. Early work on mixed-initiative clustering (Huang, 2010) demonstrated how adaptive grouping can incorporate user feedback to resolve tensions between statistical similarity and conceptual relevance. Visual interfaces like those used in Cyclone (Duman et al., 2009) and semantic interaction systems (Endert et al., 2012) enable users to reposition, merge, or split clusters, guiding the underlying models toward more meaningful structural representations. These systems treat abstraction as a dialogue between human and machine, where user intentions are surfaced through interactive spatial cues.
A complementary technique for refining abstractions is contrastive refinement—clarifying boundaries by comparing positive, negative, or near-miss examples. MOCHA (Gebreegziabher et al., 2024) illustrates this by generating counterexamples that expose the limits of a user’s current concept, prompting structural correction and alignment. Flash Fill (Gulwani, 2011) supports refinement by allowing users to correct inferred transformation rules using contrasting input-output examples. These systems show how surfacing edge cases and exceptions can accelerate convergence toward accurate and generalizable models.
Schemex builds on these traditions of mixed-initiative abstraction and contrastive refinement but targets multimodal design examples and schema-level generalization. Schemex treats structure learning as a collaborative, iterative process. It specifically supports users in moving from scattered examples to coherent, named schemas through clustering, structural abstraction, and contrastive refinement. This workflow makes schema induction more interpretable and actionable for downstream design and authoring tasks.
Figure 2. Screenshot of Schemex, an interactive, AI-powered visual workflow that helps users induce schemas from examples. Each node is interactive and features a side panel displaying detailed information, such as definitions and example mapping matrices. The workflow comprises three key stages: clustering, abstraction, and refinement through contrasting examples.
2.3. HCI Systems for Structure Discovery and Sensemaking
Schemex builds on a growing body of HCI research that supports structure discovery and sensemaking, particularly in design and multimodal content domains. In design, many systems aim to surface reusable patterns by mining common structures at scale. Prior work has emphasized documenting reusable solutions in structured schema formats (e.g., problem, context, solution) (Kruschitz and Hitz, 2010). For example, Webzeitgeist mined common UI structures from thousands of websites (Kumar et al., 2013), and other systems have identified interface design patterns across mobile apps through large-scale clustering (Alharbi and Yeh, 2015). Recent systems such as DesignWeaver (Tao et al., 2025) and Luminate (Suh et al., 2024) help users explore the design space of AI-generated variations by revealing latent design dimensions and attributes.
Structure discovery is increasingly applied to multimodal content, where visual and textual alignment reveals deeper task structures. For example, RecipeScape (Chang et al., 2018) supports this by clustering hundreds of recipes for a given dish based on procedural similarity, helping culinary professionals explore variation and identify recurring techniques at scale. Hierarchical tutorial generators (Truong et al., 2021) break down videos into stepwise instructional schemas, while VideoMix (Yang et al., 2025) aggregates how-to content into coherent workflows.
Moreover, sensemaking tools support users in organizing data, sharing insights, and externalizing evolving structures (Gero et al., 2024). Jigsaw (Stasko et al., 2007) helps users analyze document relationships through interactive visualizations, and Selenite (Liu et al., 2024) provides overviews of complex corpora using LLM assistance. To assist qualitative analysis, LLooM (Lam et al., 2024) offers a text-based approach to concept induction by enabling users to collaboratively cluster and label high-level ideas from unstructured documents. CollabCoder (Gao et al., 2024) supports team-based coding and codebook evolution. Earlier crowdsourced research further shows that externalizing schemas—whether derived by individuals or crowds—can improve ideation, analogy retrieval, and collaborative understanding (Yu et al., 2014a, b; Kittur et al., 2014).
Schemex shares the goal of supporting structure discovery through mixed-initiative workflows, targeting schema induction from multimodal content creation examples. Unlike prior systems that focus on AI-generated artifacts (e.g.,  (Suh et al., 2024; Tao et al., 2025)), Schemex emphasizes human sensemaking from real artifacts. It supports users in articulating and evolving actionable schemas. In this way, Schemex complements and extends prior systems by offering a workflow for interpretable, interactive, and grounded schema authoring.
2.4. Schemas Used in HCI Systems
Because of the generalized scaffolding schemas provide—particularly for novices—schemas are widely used in creativity support tools. Books of design patterns have influenced generations of architects (Alexander et al., 1977), software engineers (Gamma et al., 1995), and web designers (Duyne et al., 2002). In computer graphics, design patterns can help solve challenges such as laying out furniture in a room (Merrell et al., 2011), generating usable maps (Agrawala and Stolte, 2001), illustrating furniture assembly instructions (Agrawala et al., 2011), or sequencing cuts in film (Leake et al., 2017). In HCI, design patterns have been used to help novice filmmakers structure their videos (Kim et al., 2015), support human-robot interaction programming (Sauppé and Mutlu, 2014; Kahn et al., 2008), and inspire novel product ideas (Yu et al., 2014a, b). AI has also been instrumental in helping users apply abstract schemas to creative tasks such as writing stories (Mirowski et al., 2023), creating visual blends (Chilton et al., 2019, 2021; Cunha et al., 2017, 2018; Wang et al., 2023), producing trailers and teasers (Wang et al., 2024b), crafting social media videos (Wang et al., 2024a; Menon et al., 2024), and generating inclusive, persuasive emergency messages (Jit et al., 2024). In nearly all of these systems, authors have had to manually analyze examples and construct schemas by hand—a labor-intensive process. Accelerating schema induction could vastly expand access to structured creativity, enabling more people to design, communicate, and express ideas effectively.
3. Schemex System
Figure 3. Schemex Stage 1: Clustering.
Figure 4. Schemex Stage 2: Abstraction.
Figure 5. Schemex Stage 3: Contrastive Refinement.
Schemex (see Figure 2) is an interactive visual workflow that assists users in transforming a set of examples into actionable schemas through three AI-assisted stages: clustering, abstraction, and refinement via contrasting examples. If the input examples are multimodal, the system first performs a preprocessing step that converts them into structured textual representations using models such as GPT-4V and Whisper. Next, the workflow clusters the examples into distinct groups based on their latent similarities. For each cluster, the system examines the examples to derive key dimensions and dimension-specific attributes. Overall attributes of the examples within the cluster are also inferred, forming an initial schema comprising both dimensions and attributes. Finally, Schemex refines the schema through contrastive refinement: it generates outputs based on the initial schema, compares them to the original human-authored examples, and iteratively improves the schema. Throughout the workflow, each intermediate result is presented in an interactive node. Clicking a node reveals a side panel containing more detailed information, such as definitions and example mapping matrices, helping users effectively ground their schema inspection and understanding in the example set.
Schemex is implemented using the Flask/React Flow web framework. It is powered by a reasoning LLM (o3-mini-high) for reasoning and analysis tasks, and a general-purpose LLM (GPT-4o) for generation tasks—particularly in schema application. We chose o3-mini-high as our reasoning model due to its balance of accuracy and time efficiency. During early experiments, we found that it tended to generate overly profound outputs, so we reverted to GPT-4o for the generation steps. All prompts used in the system are available in Appendix A.
Next, we illustrate a typical interaction with the system through a walkthrough. In this scenario, the user wants to learn how to write an HCI paper abstract using a sample set of 15 CHI’24 best paper abstracts. The user also gathers the corresponding paper titles as input context.
3.1. Preprocessing Step for Multimodal Input
The user enters the system, which displays an example input node for specifying their goal and example set. To begin, the user enters the creation goal: “Write an HCI paper abstract”, and uploads a collection of 15 CHI’24 Best Paper abstracts along with the corresponding paper titles. The examples and corresponding inputs are stored in pairs within the system.
Although not applicable in this specific use case, the system is capable of analyzing multimodal examples. If multi-model content is detected, the system performs additional processing steps. For images, they are sent to GPT-4V to generate visual descriptions. For videos, the system extracts one frame per second as “screenshot” images and sends them to GPT-4V to generate visual descriptions and extract any available captions. The audio tracks of the videos are transcribed using Whisper. All parsed information is then combined into a structured textual file containing visual descriptions, captions (if available), and audio transcriptions.
Once the examples are processed and ready for analysis, the system generates a clustering operation node and prompts the user to click the “Run Clustering” button.
3.2. Stage 1: Clustering
Clustering (see Figure 3) identifies subclasses of examples and is crucial in preventing schema overgeneralization. When given a set of examples, it is often unclear if they share the same schema. Generalizing across structurally different examples can lead to weak and bland schemas. Therefore, we must perform clustering first to identify examples that share latent similarities with each other and then find the schema specific to each cluster.
After the user clicks the “Run Clustering” button, the system instructs the reasoning model to cluster the examples and provide reasoning. The system processes all of the examples simultaneously, producing and labeling clusters. The resulting clusters are then displayed as individual nodes on the Schemex interface.
For instance, Schemex maps the abstracts into 3 clusters: Empirical Studies (6 examples), Theoretical Contributions (4 examples) and System Design and Evaluation (5 examples).
The user can easily verify the clustering results by selecting a cluster node to view a side panel that explains the common features shared among its examples, including a matrix describing how closely each example aligns with these features. For example, Schemex identifies the common features of Theoretical Contributions papers to be: 1) Discussion of frameworks or concepts that advance understanding within the field, 2) Emphasis on implications for theory and policy, 3) Engagement with existing literature to draw new insights or reframe existing theories.
After checking the side panel, the user is satisfied with the clustering and decides to explore the Theoretical Contributions cluster in more depth. They click the ”Infer Dimensions” button on the cluster node.
3.3. Stage 2: Abstraction
Abstraction (see Figure 4) identifies structural commonalities among examples within each cluster. Unlike “features” that are used for clustering, the schemas formed during “abstraction” synthesize multiple examples into a general pattern of specific dimensions with descriptive attributes. It synthesizes multiple examples and extracts generalizable insights, a task that can often be overwhelming for individuals due to the high cognitive load involved.
Once the user clicks “Infer Dimensions” for the selected cluster, Schemex prompts the reasoning model to analyze the examples and extract the core elements or dimensions of the examples. For the Theoretical Contributions cluster, the key dimensions include: Contextual Framework , Gaps in Existing Literature, Methodological Approach, Findings and Contributions. The user can verify the generated schema by clicking on the dimension node to view further details about each dimension. This panel also includes an interactive Dimension-Example matrix that provides a visualization of how well each example conforms to each dimension. The matrix provides specific citations and justifications regarding whether each example fully, partially, or did not incorporate each dimension, enabling users to assess the system’s coherence by mapping its analysis back to the examples.
The next step is to “Infer Attributes”, including the dimension attributes, the key features within each dimension, and the overall attributes, the key descriptors of the full output. For example, for the Contextual Framework dimension, key attributes include a clearly defined population from which theoretical insights emerge or to which they apply, and the broader sociotechnical, cultural, or historical context that shapes the significance and applicability of the theory. For overall attributes, examples include a target word count range of 140–150 words and a narrative structure that presents the dimensions sequentially.
This approach distills executable guidelines for the specific cluster, ensuring that the output not only captures the key components (Dimensions) but also details what makes each component effective (Dimension-specific Attributes) and how to integrate these components into a cohesive output (Overall Attributes). At this point, the user has the first version of a usable schema.
3.4. Stage 3: Refinement via Contrasting Examples
The schema derived from Stage 2 is not perfect. For instance, while it specifies that important findings should be summarized, it does not clearly define how they should be written. To make the schema more precise and actionable, we employ an apply-and-test approach, iterating on the schema using contrastive refinement (see Figure 5).
Specifically, once the user has created an initial schema, they can click “Apply Schema” to prompt GPT-4o to generate Dimension values based on the current schema and an example’s input context (here is the paper title). This intermediate step allows users to 1) assess the accuracy of the identified dimensions and attributes before generating a final output 2) provide insight into the individual components behind the final product, deepening the user’s understanding of the system’s reasoning, as well as the task. Next, the user can generate a complete output (e.g., an HCI paper abstract) based on the generated dimension values and overall attributes. Note that the user can pick from a validation set (examples that came from the cluster was set separately and not used in the training) for schema application to check the schema effectiveness.
Schemex provides a color-coded interface that allows to user to easily compare the generated outputs with the original examples by dimension, highlighting potential improvements extracted by the reasoning model. These improvements are then used to iterate the schema. For example, the system might suggest that findings should be explicitly linked to existing literature and frameworks, rather than presented in isolation. After one iteration, the system generated a more detailed schema incorporating this improvement for the Findings dimension, such as adding the attribute “theoretical integration.” The iterated schema can then be used to generate a new set of dimension values and complete outputs. The cycle repeats until users are satisfied with the final schema. The user can keep exploring another cluster they are interested in.
Throughout their interaction with Schemex, users are involved in the process by evaluating whether the generated content meets their standards. By highlighting and citing content relevant to the AI’s reasoning decisions, Schemex encourages critical analysis of system output and constant comparison to the source material. These features support the user in their understanding of the AI and the paper abstract writing ask, and the original inputs, so they can more effectively act as a reflective evaluator.
Ultimately, users can utilize the resulting schemas as a guide to write their own HCI paper abstracts.
4. Technical Evaluation
Condition1: Condition2: Condition3: p-value p-value p-value o1-pro Schemex Initial Schemex Full (1 vs 2) (2 vs 3) (1 vs 3) D1-Cluster Quality 5.1 (0.79) 5.8 (0.72) 6.4 (0.51) 0.046 0.02 0.004 D2-Dimensions & Attributes Quality 5.0 (0.85) 5.9 (0.67) 6.8 (0.39) 0.026 0.002 0.003 D3-Coverage & Generalizability 5.0 (1.1) 5.5 (0.67) 5.8 (0.75) 0.19 0.26 0.07 D4-Example Fit 5.6 (0.9) 6.1 (0.67) 6.6 (0.51) 0.19 0.014 0.018 D5-Usefulness for Content Creation and Critique 5.3 (0.98) 6.2 (0.75) 6.7 (0.49) 0.032 0.059 0.007 D6-Alignment of Generations with Original Examples 5.1 (0.79) 5.7 (0.65) 6.4 (0.67) 0.07 0.007 0.007 D7-Overall Effectiveness of Generations 5.5 (0.52) 5.8 (0.58) 6.3 (0.65) 0.16 0.014 0.013
Table 1. Evaluation results comparing o1-pro, Schemex Initial, and Schemex Full, where means, standard deviations (in parentheses), and p-values for the paired-sample Wilcoxon tests are reported. Bolded p-values are statistically significant.
To evaluate how structural abstraction and contrastive refinement contribute to schema quality, we conducted a controlled study in which domain experts assessed schemas generated under three different conditions:
  • Condition 1: o1-pro—an advanced reasoning model capable of interpreting task descriptions, decomposing them into steps, and executing those steps to return results.
  • Condition 2: Schemex Initial—a structured workflow involving clustering, followed by abstraction into dimensions and attributes.
  • Condition 3: Schemex Full—our full system implementation with one iteration, which builds on Schemex Initial by introducing a contrastive refinement step.
We hypothesize that schemas generated under the Schemex Full condition will outperform those from the other conditions across seven dimensions: (1) cluster quality, (2) dimension and attribute quality, (3) coverage and generalizability, (4) example fit, (5) usefulness for content creation and critique, (6) alignment of schema-guided generations with original examples, and (7) overall effectiveness of schema-guided generations.
4.1. Data Preparation
We collected 6 different topics as the test suite, each containing 15 randomly picked examples. The topic span different domains of academia, business and media, and span different modalities:
  • Task 1: EE/CS faculty candidate job talk abstracts
  • Task 2: UIST teaser video
  • Task 3: LinkedIn Profile summary
  • Task 4: Pitch deck for start-ups
  • Task 5: News headline
  • Task 6: News TikToks (Newman, 2022)
Example creations are collected from trustworthy real-world resources, for example, the EE/CS faculty candidate job talk abstracts were collected from the university emails, the UIST teaser videos were downloaded from the UIST YouTube channel. All examples were processed using our multi-modal analysis pipeline to get the structured textual dataset.
To generate schema with o1-pro, we use the following prompt, which is equivalent to what we give Schemex as the task description and input. The prompt can be found in Appendix B. We ran models or pipelines in three conditions all once for each of the task. After getting all the schemas, we used GPT-4o to generate the outputs (as described in the System section “schema application” step). For each schema (of each cluster), we generated outputs for two randomly picked inputs from the cluster. Note that if a cluster just has one example, we just generated output for one input. Overall, we generated 20 schemas and 38 outputs for Condition 1, and 17 schemas and 33 outputs for Condition 2 and Condition 3, respectively.
4.2. Participants and Procedure
We invited twelve experts (two experts for each topic) to evaluate the schema and the accordingly outputs. The experts (average age 27.0, 7 female, 5 male) all have experience in the relevant domain and the specific task. For example, for news TikTok tasks, we invited journalism graduate students who had two years of experience in making news TikToks. Before the evaluation, we presented experts with examples of the schemas. We provided an evaluation rubric of the seven dimensions and guided them through it using practical examples. We requested them to read the original examples thoroughly before commencing ratings on schemas and outputs. For each of the tasks, two experts independently evaluated each dimension on a 7-point scale with justifications. When presenting the data to the experts, the condition information is removed, and the order of the condition shown to experts are counterbalanced between all the participants. The experts were compensated \$50/hour.
4.3. Evaluation Results and Findings
The interrater Agreement (within ±1 point) is 0.78, showing substantial agreement between the two experts’ judgment. This is deemed acceptable due to its highly subjective nature. We average the two experts’ scores for each schema and output. We run paired-sample Wilcoxon tests to compare the scores. See results in Table 1.
4.3.1. Comparing Schemex Initial to o1-pro: Structural abstraction significantly improves schema quality and usefulness.
Schemex Initial outperforms o1-pro in all seven dimensions, and it shows statistically significant improvements in three of the seven evaluation dimensions: D1: cluster quality (p=.046), D2: dimension and attribute quality (p=.026) and D5: usefulness of the schema (p=.032).
In terms of cluster quality, Schemex clusters are functionally exclusive and grounded in clear structural commonalities. In contrast, o1-pro’s clusters often blur together and overlap, reducing precision and limiting their usefulness. For example, in the EE/CS faculty candidate job talk case, although o1-pro-derived clusters are thematically distinct: Cluster 1: Technical or Algorithmic-focused; Cluster 2: Systems and Architecture-Focused; Cluster 3: Human and Societal Impact-Focused. Example 10 (which is about supporting user-centered analytical interface at scale) was assigned to Cluster 2 due to its system focus, but can also be placed to Cluster 3 as it also highlights the “human” perspective. While Schemex-derived clusters have a clear structural divide which each example can be cleared mapped to: Cluster 1: Sequentially Ordered Roadmap Abstracts and Cluster 2: Integrated Narrative Abstracts.
In terms of dimension and attribute quality (D1), Schemex offers more granular, example-set-specific attributes that closely reflect structural patterns in the data. o1-pro, by comparison, relies on broad, vague guidelines that lack depth and fail to capture nuances. For example, in the LinkedIn profile summary case (D2), while o1-pro notes that they “often include structured bullet points” or “end with an invitation to subscribe,” Schemex identifies precise narrative moves like a “pivotal challenge or turning point,” to detail career evolution—offering clearer insights into how personal storytelling and structural choices actually function in the examples.
In terms of the usefulness of schema (D5), Schemex provides practical, actionable guidance that helps users emulate gold examples effectively. o1-pro is less descriptive and offers limited support for creation or critique. For example, for the UIST demo system walkthrough cluster, while o1-pro mentions key elements like on-screen annotations, Schemex goes further by detailing how annotations are layered to guide user understanding—making it more actionable for creators aiming to emulate complex interactive walkthroughs like those in the example set.
4.3.2. Comparing Schemex Full to Schemex Initial: Contrastive refinement unlocks additional gains for example fit and generation quality.
Schemex Full goes a step further, yielding statistically significant improvements in 5 out of 7 dimensions compared to Schemex Initial: D1: cluster quality (p=.02), D2: dimension and attribute quality (p=.002), D4: Example fit (p=.014), D6: Alignment of Generations with Original Example (p=.007) and D7: Overall Effectiveness of Generations (p=.014).
As Schemex Full went through another cycle of contrastive analysis and abstraction, it is not surprising that it can generate higher-quality schema (D1 and D2). In terms of example fit (D4), through contrastive refinement, Schemex Full enables finer-grained alignment between schema attributes and examples, capturing subtle structural cues that Schemex Initial often missed. For example, for the news headline case, the initial version emphasizes using active verbs and focusing on a single event. The full version adds guidance for compound action handling—such as using “and,” “as,” or “after” to connect multiple developments—which better reflects several gold examples that combine actions and consequences in a single headline. This also explains why schemas derived by Schemex Full can generate outputs that better align with the original examples (D6) and serve the creation purpose more effectively (D7), as it captures more nuanced rules from good examples for AI to follow.
Participant Example set Number of Examples
P1 Personal creative writing 10
P2 NBA team’s Instagram posts 13
P3 Iambic Pentameter 15
P4 Subheaders from NPR Articles 14
P5 Tech Career LinkedIn Bios 8
P6 HCI PhD SoP First Paragraph 12
P7 Fandom Wiki for Harry Potter 10
P8 Hopecore TikTok 8
Table 2. Participants and the example sets they choose.
4.3.3. Limitations of Contrastive Refinement: Coverage and Generalizability.
Interestingly, D3—Coverage—is the only dimension where Schemex Iterated did not show a statistically significant improvement over o1-pro (p = 0.07). This dimension evaluates how well a schema accounts for the range of structural patterns that might appear in unseen or future examples, without overfitting to the initial sample set. This result suggests that while contrastive refinement improves precision, depth, and stylistic alignment, it may not inherently expand the generalizability of the schema. In fact, by refining based on specific gaps relative to a gold example, the iteration might even bias the schema further toward that particular instance—narrowing rather than broadening its scope. By contrast, o1-pro, as a general-purpose reasoning model, may maintain broader coverage.
For example, for the pitch deck test suite, one limitation of the Schemex full-version schema is that it places greater emphasis on specific storytelling patterns—such as step-by-step, roadmap-style narratives—which were prominent in the example set. While this leads to more precise alignment with those patterns, it may not generalize as well to pitches that follow less conventional structures. For instance, a founder who presents their mission, product, and traction in a fluid, vision-led narrative—without using explicit sequencing—might be harder to fully capture within the refined schema. The core logic is still there, but the schema may not recognize it as strongly. In contrast, the o1-pro schema is more focused on identifying the underlying reasoning—problem, solution, and evidence—regardless of how the story is told. This gives it more flexibility in handling diverse and novel formats, even when they diverge from patterns seen in the example set.
Overall, our findings demonstrate that structural abstraction (Schemex Initial) significantly improves schema quality and practical utility over a strong baseline (o1-pro), particularly in cluster coherence, attribute granularity, and actionable guidance. Building on this, the addition of contrastive refinement (Schemex Full) yields further gains in alignment with examples and generation effectiveness by capturing subtler structural nuances. However, this added precision may come at the cost of generalizability, suggesting a trade-off between depth and breadth in schema induction.
5. User Study
Figure 6. Participant ratings over five research questions for baseline vs. Schemex.
To understand how Schemex assists users in schema induction, we conducted a within-subjects study with eight participants, comparing their interaction with Schemex to a baseline of ChatGPT (o1-pro). We chose o1-pro as a strong language model baseline that excels in complex problem-solving, enabling us to isolate the specific benefits of Schemex for schema induction tasks.
5.1. Protocol
We recruited eight participants (average age=26.6, four female, four male) for this study through a university mailing list and word-of-mouth. All participants reported being familiar or very familiar with ChatGPT. Each participant selected an example set of their choice. The example sets spanned diverse areas and modalities with 8-15 examples per set (see Table 2).
The participant was then given two tasks, (Task A) inducing a satisfactory schema using Schemex, and (Task B) inducing a satisfactory schema using o1-pro for the same example set. Half began with Task A, and half with Task B, with a brief demonstration before the Schemex task. Each task had a time limit of 30 minutes. After each task, participants were interviewed regarding their experience with the corresponding system. In these interviews, users rated questions (see questions in Section 5.2) on a 1-7 scale and provided reasons for their ratings. Each study session lasted approximately 80 minutes, with participants compensated \$25/hour. We run paired-sample Wilcoxon tests to compare the ratings.
5.2. Findings
Schemex significantly outperformed ChatGPT (o1-pro) in 4 out of 5 areas that were analyzed, including structural insight, novel reflections, confidence, and visual presentation. In terms of clarity, both Schemex and ChatGPT scored similarly. To assess performance across these areas, users rated the five research questions on a 1–7 scale. See results in Figure 6.
  • RQ1 - Schema Clarity: How clear and understandable was the schema provided by the system?
  • RQ2 - Structural Insight: How much new insight did you gain into the structure or norms of this genre?
  • RQ3 - Novel Reflection: Did interacting with the system help reflect or notice things you otherwise wouldn’t have?
  • RQ4 - Confidence: After using this system, how confident do you feel about the schema?
  • RQ5 - Visual Presentation: Do you think the UI/visual presentation helps you learn or enhance your understanding?
Qualitative analysis of user interviews suggests that the high ratings for Schemex stemmed not just from the outputs, but from the interactive workflow that guided users through the schema creation process. These interactions deepened users’ understanding of their original material and increased trust in the system’s output. We discuss specific quantitative and qualitative findings for each research question below:
5.2.1. RQ1 - Schema Clarity
Schemex provided a similarly clear schema (5.56/7) as ChatGPT (5.69), p=.680.
While both systems were provided comparable scores, the clarity of the two system’s final schemas came from very different areas. P8 (Hopecore TikToks) scored both Schemex and ChatGPT 6/7. She explains that while ChatGPT “was pretty readable [and] stayed kinda high-level,” Schemex had “a lot going on, but it was laid out super visually - it was easy to follow.” Overall, users found that GPT’s simple, text-based output was comparably clear to Schemex’s structured, information-rich output.
5.2.2. RQ2 - Structural Insight
Schemex provided significantly greater structural insight (5.56/7) than ChatGPT (2.25/7), p = .004.
Schemex was able to capture deeper semantic structures than the baseline. P1, who analyzed diverse samples of her writing, rated Schemex a 7/7, citing its identification of common thematic threads in her work, despite their structural differences: “it found the implicit connections and anxiety around time.” The ability to gradually break down the inferred commonalities into specific examples through interaction with Schemex encouraged intellectual engagement. P2 continued, “[Schemex] tells you - ‘this is a thing you do and here’s the precise way it manifests.’ It makes me reflect.”
Additionally, Schemex outputs were more detailed, providing more precise information than the o1-pro model. P3, who was analyzing various examples of iambic pentameter, rated Schemex 7/7. His final schema provided specific instructions regarding meter, language style, and how to form the rhetorical questions characteristic of classical poetry in the genre.
Schemex’s depth and detail made it the participants’ preference for tasks requiring a personal understanding of the material. For example, P3 (Iambic Pentameter) noted that Schemex would be very helpful if he was trying to teach a course on the subject. P5 (LinkedIn Bio), who provided Schemex with a score of 4 and the baseline with a score of 2, echoed, “If I was trying to understand what a LinkedIn bio is, I would go to [Schemex].”
Figure 7. Schemas P4 induced with o1-pro (left) and Schemex (right) for the example set of NPR article subheaders. While ChatGPT gave a generic outcome, Schemex provides clusters, dimensions, and attributes that P4 found insightful and specific to the given example set. Inside the dotted box is one output from Schemex and the according gold example.
5.2.3. RQ3 - Novel Reflection
Schemex encouraged a higher level of reflection and discovery (6.06/7) compared to ChatGPT (2.44/7), p = .004.
Interacting with Schemex sparked new insights participants had not noticed on their own. Many users initially believed their example sets were straightforward. However, Schemex’s intermediate steps, particularly clustering, helped surface unexpected distinctions. P4, analyzing NPR subheaders, discovered meaningful categories like “Direct Factual Summaries” and “Composite and Contrasting Narratives.” See Figure 7 for the detailed schema. He scored Schemex 5/7, explaining: “It helps you understand something that seems like ‘the same thing’ [and find] there’s different types of them.”
Furthermore, exploration options provided by Schemex provided avenues for refinement they had not considered. For example, when using ChatGPT, P2 (Instagram Posts) had difficulty determining how to improve its outputs: “You have to know what to ask it in order to get what you want. It’s a little bit more frustrating.” He scored ChatGPT 2/7. P1 explains how the options provided by Schemex (scored 7/7) remedy this problem, citing the suggestions provided in the comparison step: “Schemex helps you with how you should iterate and provide you with an understanding of all of the opportunities.” He compares this to the experience using ChatGPT (scored 2/7): “I know what needs to change, but I don’t know what to change to. [I have to] act in a very corrective way.”
5.2.4. RQ4 - Confidence
Participants expressed greater confidence in schemas produced by Schemex (5.75/7) than ChatGPT (3.88/7), p = .011.
Schemex’s layered, interactive workflow fostered faith and confidence in outputs. P3 shared: “Because [at each level] I could see the added context, it was a lot easier to trust - which wasn’t the case with ChatGPT. If I had used ChatGPT second, I would have given [ChatGPT] lower scores.”
Schemex further instilled confidence in users by providing them with accessible and efficient evaluation features. P2 emphasized the example matrix as a confidence-boosting element: “It was just an easy way to sort through [the information] … and not have to go back and figure out each one.” P5 added: “I liked the self-reflection of the AI. It made me feel more confident in what it was telling me.”
Schemex developed users’ trust by grounding abstract representations in real data. P1, when discussing the limitations of AI systems, noted, “I love that it helped balance the abstractness…I don’t think an AI-empowered interface could do without examples.” In contrast, the same participant highlighted a key shortcoming of ChatGPT, stating, “the way ChatGPT engaged with the examples was generally summarizing.” P5 similarly emphasized the role of concrete examples in fostering trust, explaining their 7/7 rating by saying, “I think it also goes to the trust in the system. Examples are the most helpful for me.”
Ultimately, users were confident in the precision, depth, and actionability of the Schemex schemas. P3 noted: “It was just so much more descriptive and aligned with what I hope I would take away if I had a lot of time spent studying this.” P5 echoed this confidence regarding her Schemex output: “I would have been able to write a LinkedIn bio [from it].”
5.2.5. RQ5 - Visual Presentation
Schemex’s visual interface (6.0/7) was rated significantly higher than ChatGPT’s (2.88/7), p = .008.
Schemex’s “tree-like” layout more closely matched participants’ conceptual model of a schema than linear text. P1 explains why the Schemex interface better represents schemas when compared to ChatGPT: “Schemas are abstractions. They’re less linear. It’s harder to conceptualize and bucket things in pure text.” P3 agreed that Schemex allowed for improved information processing: “Schemex was by far the most granular at visualizing and demonstrating all the specific details.”
Schemex’s interface enhanced user understanding by allowing users to “drill down” from abstract groups to general insights to specific details. The “overview first, zoom and filter, then details on demand” mantra, or Visual Information-Seeking Mantra, suggests presenting data effectively by first offering a broad overview, then allowing users to explore specific details through interactive tools. P1 enjoyed how this structure allowed her to gradually break down patterns in her writing. She explains, “It started off with broad themes, [then] an integrated framework, and [then] how you capture ‘vivid imagery’ in your writing specifically with an operational definition.” P2 also found that this step-by-step breakdown made the logic of the system easier to follow, stating, “I like how it drilled down, it gave me more insight into the AI and how it determined the less obvious stuff.”
However, not all participants preferred the Schemex interface. P5 scored Schemex 3/7 and ChatGPT 4/7, expressing her preference for more condensed, linear representations of information: “I am the kind of person who likes to see what I am focusing on… [Schemex] was a little too much for me.” Despite this, the majority of participants found the visualization to significantly support their understanding.
6. Discussion
6.1. Supporting Schema Discovery through Interactive Structural Abstraction
Schemex scaffolds interactive structural abstraction through AI-assisted reasoning and visualizations. In our study, participants consistently preferred the interactive visual workflow of Schemex over chat-based interfaces and reported gaining deeper insights. This preference is not surprising in retrospect: learning benefits from active engagement. Cognitive theories such as constructivism (Fosnot, 2013) emphasize that people learn more effectively when they can manipulate, reflect on, and iterate over ideas. In contrast, chat interfaces like GPT-style models often promote passive interaction—users tend to skim for answers rather than engage in deeper exploration. By making structure visible and refinable, Schemex encourages users to think more critically and deliberately, leading to more meaningful learning and insight.
Discovering underlying structure from examples is rarely a linear task. Like other sensemaking processes, schema induction involves generating hypotheses, testing them against data, and refining them based on feedback. Users often need to zoom in on fine-grained patterns and zoom out to adjust the overall schema. Visual workflows like Schemex make this iterative process manageable by scaffolding multiple levels of abstraction—from clusters to dimensions to example-level attributes. This mirrors how experts build understanding over time and enables novices to follow a similar path through guided exploration.
The effectiveness of interactive schema discovery also depends on the capabilities of the underlying AI. Large language models—especially those optimized for reasoning—have recently become capable of meaningful structural analysis: clustering examples, identifying latent dimensions, and proposing abstractions that align with human interpretations. Schemex leverages these capabilities through a guided, multi-stage process: the AI suggests patterns, but always in a way that users can inspect and refine. This makes the system a collaborative partner rather than a black box, aligning well with best practices in human-AI interaction.
While Schemex was designed and evaluated for schema induction in creative and communicative tasks, the underlying approach is applicable to a wide range of domains. Many real-world problems involve extracting structure from complex, nonlinear data—for example, mapping workflows in organizational process mining, modeling user flows on websites, or tracing state transitions in software systems. These domains all require iterative exploration and abstraction to surface useful insights. The interactive structural abstraction model supported by Schemex provides a general framework for approaching such problems, suggesting promising directions for future applications.
6.2. Contrastive Refinement as a Mechanism for Evolving Schema Understanding
People often learn by comparing examples—noticing variations, identifying what holds constant, and refining their understanding through contrast. Schemex builds on this principle through contrastive refinement, where it compares AI-generated outputs to real examples to evaluate and adjust the underlying schema. This process helps recognize mismatches, clarify vague structures, and iteratively improve schema quality.
Contrastive methods are already central in AI and machine learning, particularly in contrastive learning approaches, where models improve by distinguishing between similar and dissimilar instances. In Schemex, contrast serves a different but related purpose: it helps refine abstract representations through guided comparison. This aligns with cognitive theories of structural alignment (Markman and Gentner, 1993), which emphasize that abstraction is often clearest when comparing systematically varied cases.
Importantly, schemas are rarely static. As genres shift or new formats emerge, existing schemas may no longer fit. For instance, an HCI researcher familiar with traditional system papers may encounter first-person HCI work and struggle to interpret its structure using previous templates. Contrastive refinement supports this kind of reflection by highlighting where existing schemas fall short and how they might adapt. It enables users to remain flexible and responsive as domains evolve.
Future work could extend this idea toward longer-term schema tracking. Schemas may shift over time within individuals or communities: evolving, merging, or splitting as practices change. Supporting experts in monitoring these shifts could help them stay current, inform pedagogical tools, or uncover emerging patterns in fields. Contrastive refinement offers a foundation for such reflective exploration by anchoring schema development in real, observable differences.
6.3. Limitations and Future Work
Schemex currently supports multi-modal example analysis by converting images, videos, or audio into structured textual descriptions to enable LLM-based reasoning. While this approach allows schema induction across diverse content types, it limits the evaluation and application of schemas to a primarily textual modality. Future work could explore fully multi-modal interaction—for example, aligning schema elements with visual or interactive outputs—to help users better assess how schemas function across different media.
Schemex also includes a basic schema application workflow: given a user input and a derived schema, the AI generates content by filling in values for each dimension. This feature helps users test the schema in context and refine it through contrastive evaluation. However, supporting real-world schema application more effectively presents open challenges. What kinds of inputs should users provide? Given any schema, how can we scaffold users through the process of applying it meaningfully? What forms of intervention or guidance should the system offer? Future research could explore how to tailor this workflow to different content domains, input types, and levels of user expertise.
Finally, the potential misuse of Schemex raises ethical concerns, such as scraping others’ examples or imitating stylistic patterns without proper attribution (Porquet et al., 2024). Addressing these issues will be essential as the system moves toward broader deployment. Future versions of Schemex could incorporate mechanisms for attribution tracking, usage transparency, and community norms to encourage responsible use.
7. Conclusion
We introduced Schemex, an interactive system that supports schema induction through AI-assisted structural abstraction and contrastive refinement. By grounding abstract patterns in real examples and enabling iterative comparison, Schemex helps users uncover, evaluate, and refine the underlying structure of creative artifacts. Our study shows that Schemex improves user insight, confidence, and schema quality compared to a reasoning AI model baseline. This work highlights the potential of interactive, example-driven workflows to scaffold complex reasoning tasks, and opens new directions for supporting structure discovery across domains.
Appendix A Prompts used in Schemex
A.1. Get clusters
I’m learning how to {content_type}.Analyze the following examples to identify clustersbased on STRUCTURAL and RHETORICAL patterns.Focus on how the content is constructed andpresented—not just what it’s about.For textual examples, examine rhetorical organization.For multimodal examples, consider how differentmodalities interact structurally and rhetorically.For each cluster:Name the structural approach (not just the topic)List 2–3 shared structural featuresInclude the example IDs in that clusterClustering Rules:Assign every example to exactly one clusterUse all examples—no omissions or duplicatesGroup based on structure, not content themes or IDorderResponse Format:Cluster 1: [Cluster Name]Common Features:- [Feature 1]- [Feature 2]Examples:- Example [ID]- Example [ID]...Total number of examples: [Number]Cluster 2: [Cluster Name]...Examples to analyze: {examples}{input_context}
A.2. Get the feature-example matrix
I’m learning how to {content_type}.Help me analyze how each example in the cluster{cluster_name} demonstrates the identified commonfeatures.Common Features: {common_features}Examples: {examples}Your Task:For each example:Indicate whether it demonstrates each feature:"Yes" = Clearly demonstrates the feature"Partial" = Feature is present but limited, implied,or modified"No" = Feature is not present or is inconsistent withthe definitionInclude:A brief explanationA direct snippet from the example (if marked "Yes" or"Partial")Format Requirements:Return your output as a valid JSON object with thefollowing structure:{  "mapping": [    {      "example_index": "EXAMPLE_ID",      "example_snippet": "First 50 characters of the      example...",      "feature_mapping": [        {          "feature": "Feature 1",          "feature_id": "F1",          "applies": "Yes/No/Partial",          "explanation": "Brief explanation of          classification",          "snippet": "Verbatim quote from example that          demonstrates the feature"        },        ...      ]    },    ...  ]}
A.3. Infer dimensions and get the dimension-example matrix
I’m learning how to {user_goal}.Analyze the following examples from the cluster:{cluster_name}Examples: {examples_str}{input_context}Your Task:Help me infer the structural commonalities thatdefine this cluster. Focus on identifying the keycontent dimensions that shape the narrative andcomposition.Important Constraints — AVOID OVERFITTING:Only identify dimensions shared across multipleexamplesFocus on cluster-wide structural patterns, notindividual quirksBroaden dimensions only if needed to make them generalenough for the clusterFor Each Dimension:Give it a clear, descriptive nameBriefly explain what the dimension capturesInclude verbatim snippets from examples thatdemonstrate itFor Each Example:For every dimension, show:Whether it applies: Yes, Partial, or NoA short explanation of whyA verbatim snippet from the example (if marked "Yes")If no snippet can be found, use "Partial" or "No"Snippet Rules:Only use exact text from the examples—no paraphrasingor made-up textIf a snippet can’t be found, say so—never fabricateoneApplies Judgment:Yes = Strong, specific snippet availablePartial = Present, but limited or implicitNo = Absent or not applicableOutput Format:Return your response as a valid JSON object.{  "dimensions": [    {      "name": "Dimension name",      "description": "Brief explanation",      "examples": [        {          "example_id": "1",          "snippet": "Verbatim snippet from Example 1"        },        ...      ]    },    ...  ],  "example_mappings": [    {      "example_id": "1",      "dimension_applications": [        {          "dimension": "Dimension name",          "applies": "Yes/No/Partial",          "explanation": "Brief explanation",          "snippet": "Exact text from this example"        },        ...      ]    },    ...  ]}
A.4. Infer dimension-specific attributes and get the attribute-example matrix
I’m learning how to {user_goal}.Please analyze the following examples from thecluster: {cluster_name}Examples: {examples_full_text}{input_context}PART 1: IDENTIFY DIMENSION ATTRIBUTESYou’ve already identified the following dimensions:Dimensions: {dimensions_text}Your task now is to define key attributes under eachdimension, based on patterns that are consistentacross the examples in this cluster.Use only the following example IDs: Example IDs:{example_ids_text}Attribute Requirements:Each attribute must appear in at least 50% of theexamplesKeep attributes broad enough to apply across examples,but still specific and observableExclude attributes that are too narrow, ambiguous, orinconsistently implementedEvaluate every example against every attribute—noomissionsFor Each Dimension, Provide:Detailed Attributes: Clear, concrete, 1-sentencedescriptions of each attributeConcise Summaries: 1–2 word phrases summarizing eachattribute (in the same order)PART 2: ATTRIBUTE IMPLEMENTATION IN EXAMPLESFor every attribute in Part 1, assess how each exampleimplements it.Use the following classifications:"YES" — Clearly and fully present (must include adirect quote)"PARTIAL" — Present but limited, modified, or implicit"NO" — Not present or structurally inconsistent withthe attributeImplementation Guidelines:Always use the actual example IDs from the datasetProvide a verbatim quote from the example if theclassification is “YES” (never paraphrase)Include an explanation for every example, even if “NO”Never fabricate text—if you can’t find a quote, use"PARTIAL" or "NO"Evaluate every example for every attribute, even ifit’s “NO”Critical Format Requirements (Strict JSON Only):{  "dimensions": {    "Dimension Name": {      "detailed": ["Detailed attribute 1", "Detailed      attribute 2", ...],      "concise": ["Summary 1", "Summary 2", ...]    }  },  "attributes_examples": {    "Dimension Name": {      "Detailed attribute 1": [        {          "example_id": "ACTUAL_ID_1",          "quote": "Exact quote from example",          "explanation": "How this quote demonstrates          the attribute",          "classification": "YES"        },        {          "example_id": "ACTUAL_ID_2",          "quote": "",          "explanation": "Why this example does not          include this attribute",          "classification": "NO"        }      ],      ...    }  }}
A.5. Infer overall attributes and get the attribute-example matrix
I’m learning how to {user_goal}.Please analyze the following examples to identifyoverall structural patterns that are consistentacross most of them.Examples: {examples_full_text}{input_context}Use only these example IDs in your analysis: ExampleIDs: {example_ids_text}Goal: Identify Overall AttributesYour task is to identify broad structural orcompositional attributes that appear across themajority of examples, regardless of individualdimensions.Look for consistent patterns in areas like:Length – Word/sentence count range that appears inmost examplesFormat – Layout elements (e.g. paragraph count,headers, visual markers)Tone – Overall voice, formality, or affective stanceOrganization – Common sequence or structure acrossexamplesOther global traits – Any other recurring patternsseen in most examplesYou must include at least one attribute about lengthor format, and one about organization.Attribute GuidelinesEach attribute must appear in at least 50% of theexamplesKeep attributes broad enough to apply across manyexamplesGround attributes in observable patterns (notinterpretations)For Each Attribute:Provide a detailed description: a one-sentenceexplanation of the observable patternProvide a concise summary: a 1–2 word label for eachattributeThe number of concise summaries must match the numberof detailed attributes.Attribute Application per ExampleFor each attribute, evaluate every example:YES = Attribute is clearly and fully present (mustinclude direct quote)PARTIAL = Attribute is present but limited, modified,or implicitNO = Attribute is absent or structurally inconsistentFor every example, include:Verbatim quote from the example (if applicable)A brief explanation of your classificationClassification: "YES", "PARTIAL", or "NO"STRICT FORMAT (DO NOT DEVIATE):{  "overall_attributes": {    "detailed": ["Detailed attribute 1", "Detailed    attribute 2", "Detailed attribute 3"],    "concise": ["Concise1", "Concise2", "Concise3"]  },  "overall_attributes_examples": {    "Detailed attribute 1": [      {        "example_id": "ACTUAL_ID_1",        "quote": "Exact quote from example",        "explanation": "Why this example demonstrates        the attribute",        "classification": "YES"      },      {        "example_id": "ACTUAL_ID_2",        "quote": "",        "explanation": "Why this example does not        include the attribute",        "classification": "NO"      },      ...    ],    "Detailed attribute 2": [...],    "Detailed attribute 3": [...]  }}
A.6. Get dimension values
I am trying to {current_user_goal}.Given the following input: {input_text}I need you to generate only one component of theoutput—not the full output.Component to generate: {dim_name} – {dim_description}Component Requirements: {attributes_text}Important:Focus only on generating the specified component.Do not include other parts of the output.Follow the given requirements closely.
A.7. Get contrasting examples
I am trying to {current_user_goal}.Given this input: {input_text}And these dimension values: {dimensions_text}Help me {current_user_goal} by generating a completeoutput.Make sure to:Integrate the dimension values effectivelySatisfy the following overall requirements:{overall_arrtibutes}Output the final content directly.
A.8. Get comparative analysis
Your task is to analyze the gap between a generatedoutput and a gold/reference example, in order toimprove the schema itself.Goal:Your focus is on identifying what the schema ismissing that would help the generated outputbetter align with the gold example in moremeaningful and sophisticated ways.Inputs:Schema: {schema_text}Dimension Values: {dimension_values_text}Generated Output: {generated_output}Gold Example (Reference): {gold_example}For Each Dimension:Your analysis should address:Patterns or qualities in the gold example that are notcaptured in the current schemaGaps between the generated output and the gold examplethat a stronger schema could help bridgeDeeper or more nuanced attributes that could make theschema more generative, precise, and aligned withhigh-quality outputsThink About:What deeper structural or rhetorical qualities existin the gold example that is missing from theschema?Where is the schema too shallow, vague, or rigid?How can the schema be refined or extended withoutoverfitting?For Each Dimension, Suggest 2–3 Improvements, Usingthese Tags:[ADD] — Introduce a new attribute or pattern[DEEPEN] — Make an existing attribute moresophisticated or layered[REFINE] — Clarify or better define an existingattribute[RESTRUCTURE] — Reorganize or reconceptualize thedimensionAlso, include a section for Overall(cross-dimensional insights).Output Format (STRICT):Return your response as valid JSON using the followingstructure.{  "dimension_analysis": {    "Dimension Name 1": {      "analysis": "Analysis of deeper patterns in this      dimension that aren’t captured by the      schema",      "improvements": [        "[ADD] Description of a new attribute",        "[DEEPEN] Suggestion for making an existing        attribute more sophisticated",        "[REFINE] Clarify an attribute definition",        "[RESTRUCTURE] Suggest restructuring if needed"      ]    },    "Dimension Name 2": {      "analysis": "Analysis of missing or underdefined      qualities relevant to this dimension",      "improvements": [        "[ADD] ...",        "[DEEPEN] ..."      ]    },    "Overall": {      "analysis": "Insights about patterns or      improvements needed across dimensions",      "improvements": [        "[ADD] Cross-dimensional schema improvement",        "[RESTRUCTURE] Suggest schema-wide        restructuring"      ]    }  }}
A.9. Get color coding
Your task is to create a detailed, structuredcomparison between two texts using schemadimensions to guide your analysis.TEXTS TO COMPARE:TEXT 1 (Generated Output): {generated_output}TEXT 2 (Gold Example): {gold_example}SCHEMA DIMENSIONS: {schema_text}TASK:Segment both texts and align corresponding partsside-by-side. For each segment:Indicate which schema dimension (if any) the segmentprimarily reflectsProvide a brief annotation that explains the quality,relevance, or difference of this segment comparedto its counterpartInclude the start and end character indices from theoriginal source textSet "dimension": null if the segment doesn’t clearlyrelate to any dimensionRate the importance of this segment as "high","medium", or "low" based on its role in conveyingschema-aligned meaning or structureCRITICAL REQUIREMENTS:Cover the full content of both texts—no omissionsallowedDo not duplicate content across segments—eachcharacter should appear in exactly one segmentAccuracy of start_index and end_index isessential—these must correspond exactly to thecharacters in each original textUse actual JSON null (not the string "null") when asegment doesn’t map to any dimensionOUTPUT FORMAT (Strict):{  "segments": [    {      "id": "segment_1",      "source": "generated" | "gold",      "text": "text of the segment",      "start_index": 0,      "end_index": 42,      "dimension": "Dimension Name" | null,      "annotation": "Brief analysis of the segment",      "importance": "high" | "medium" | "low"    },    ...  ],  "dimension_analysis": {    "Dimension Name 1": "Summary analysis of how this    dimension compares across the two texts",    "Dimension Name 2": "Additional dimension    summary...",    ...  }}
A.10. Get iterated schema
You are helping improve a schema for AI-guidedgeneration.User Goal: {user_goal}Context: {context_text}The original schema includes:Dimensions (each with detailed + concise attributes)Overall attributes (for general output quality)Your Task:Use the suggestions below to iterate and improve theschema by:Revising dimension attributes for clarity, depth, orgeneralityAdding new attributes where necessaryImproving overall attributes to enhance global outputqualityRemoving or simplifying overly specific attributesthat may cause overfittingUse These Inputs:Suggested Improvements: {all_suggested_improvements}Original Schema: {original_schema}Critical Format Requirements:For every dimension:"detailed" and "concise" arrays must existEvery "detailed" attribute must have a corresponding"concise" one, and vice versaFollow the exact same JSON structure as the originalschemaFinal Output:Return a valid JSON object with the same structure asthe original schema, reflecting all suggestedimprovements.
Appendix B Prompts used for o1-pro in technical evaluation
I want to learn how to {content_type}.Here are {number_of_examples} examples. Can you helpme infer the schema (structural commonalities)among the examples? So that I can use them laterto create my own. You can give me multiple schemasif there are subgroups amongst the examples - eachsubgroup has one different schema.For each schema, please outline the key dimensions(characteristic features), their correspondingattributes (descriptors of the dimension), andoverall attributes (general descriptors of thestructure and content of the final output).Give me the answer in the format: for each schema,give it a descriptive name, list correspondingexamples, overall attributes, dimensions, andattributes for dimension.Examples to analyze: {examples}{input_context}